### v1

# Hardware-Aware Quantum Kernel Design Based on Graph Neural Networks

Abstract
Quantum kernel methods have emerged as a promising direction in quantum machine learning (QML), offering a principled way to map classical data into high-dimensional quantum Hilbert spaces. While conceptually powerful, designing effective quantum kernels that adapt to both the target task and the constraints of near-term quantum hardware remains a nontrivial challenge. In this work, we propose HaQGNN, a hardware-aware quantum kernel design framework that integrates quantum device topology, noise characteristics, and graph neural networks (GNNs) to evaluate and select task-relevant quantum circuits. By predicting surrogate metrics related to fidelity and kernel performance, HaQGNN enables efficient circuit screening at scale. Feature selection is further incorporated to improve compatibility with limited-qubit systems and mitigate kernel degradation. Extensive experiments on three benchmark datasets, Credit Card, MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms existing quantum kernel baselines in terms of classification accuracy. Our work highlights the potential of learning-based and hardware-aware strategies for advancing practical quantum kernel design on NISQ devices.
preprint: APS/123-QED
I Introduction
Quantum computing presents a compelling prospect for revolutionizing machine learning tasks by exploiting the fundamental properties of quantum mechanics, such as superposition, entanglement, and high-dimensional Hilbert spaces [1, 2]. Quantum machine learning (QML), as a important class of quantum algorithms tailored for the Noisy Intermediate-Scale Quantum (NISQ) era [3], has demonstrated practical applicability in diverse real-world tasks, including classification [4, 5, 6, 7, 8], generative modeling [9, 10, 11, 12], and learning physical systems [13, 14]. Among various QML approaches, quantum kernel methods have attracted increasing attention due to their solid theoretical foundation and interpretability [15, 16, 17]. By embedding classical data into quantum states through parameterized quantum circuits, quantum kernels define implicit feature maps capable of capturing complex nonlinear relationships beyond classical means.
Despite recent demonstrations of quantum kernel learning on real quantum devices, realizing practical quantum advantage remains challenging, especially under the constraints of Noisy Intermediate-Scale Quantum (NISQ) devices [18, 19, 20]. In particular, unsuitable quantum circuit designs may lead to vanishing kernel similarities [21], reduced expressivity, and degraded generalization [22, 23], severely limiting the performance of quantum kernels. These drawbacks are exacerbated when the data dimensionality is high or when the circuit depth and qubit count exceed the noise tolerance of near-term devices. To mitigate these limitations, several works have investigated circuit-aware and data-aware quantum kernel optimization. Some methods focus on tuning variational parameters in fixed circuit layouts [24, 25, 26], while others explore structural search via evolutionary or Bayesian algorithms [27, 28, 29]. However, most existing approaches either ignore hardware constraints such as qubit connectivity and gate noise, or suffer from excessive computational costs in searching large quantum circuit spaces.
Figure 1: An overview of handware-aware quantum kernel design.
To address these challenges, we propose a hardware-aware quantum kernel design framework based on Graph Neural Networks (HaQGNN), which efficiently integrates quantum hardware topology and device noise characteristics into the design of quantum feature maps, as illustrated in Fig. 1. The use of GNNs builds upon our prior work [30], where graph-based models were employed to predict specific properties of quantum circuits with high accuracy. Rather than formulating quantum kernel design as a joint optimization over circuit structure and parameters, our approach decouples the process by first generating hardware-aware candidate circuits and then employing graph neural networks to evaluate their potential performance. Leveraging the graph structure of quantum circuits, we employ GNNs to predict two surrogate metrics: the ”Probability of Successful Trials (PST)” [31], which reflects circuit fidelity under hardware noise, and kernel-target alignment (KTA), which correlates with downstream classification accuracy.
The HaQGNN framework consists of four steps: (1) candidate circuit generation based on hardware topology and native gate sets; (2) PST prediction via GNNs-1 to filter out low-fidelity circuits; (3) KTA prediction via GNNs-2 to select high-performing kernels; and (4) parameter optimization of the top-$k$ circuits for final task-specific performance. Furthermore, we incorporate feature selection to reduce input dimensionality, mitigating vanishing similarity issues and enhancing compatibility with NISQ devices. To the best of our knowledge, this is the first work to propose a GNN-based, hardware-aware framework that evaluates and selects quantum kernel circuits by jointly considering hardware topology, noise characteristics, and circuit structure, followed by task-specific parameter optimization. Extensive simulations on IBM quantum devices demonstrate the superiority of HaQGNN over several state-of-the-art quantum kernel baselines across multiple classification tasks and device configurations.
Our proposed HaQGNN framework offers several distinctive advantages that collectively enable efficient quantum kernel design for NISQ devices:
  • $\bullet$
    Hardware-awareness: HaQGNN explicitly incorporates real-device topology and noise information during both circuit generation and performance prediction, enhancing compatibility and circuit fidelity across different quantum hardware backends.
  • $\bullet$
    Two-stage GNN-based performance estimation: By employing two lightweight graph neural networks (GNNs-1 and GNNs-2), the framework accurately predicts fidelity (PST) and classification potential (KTA) of circuits with minimal simulation cost, achieving speedups of over 380× and 58000× compared to traditional metrics.
  • $\bullet$
    Noise-robust candidate selection: By filtering out low-fidelity circuits early in the pipeline, HaQGNN effectively mitigates the impact of gate errors and decoherence, leading to more reliable kernel performance on noisy devices.
  • $\bullet$
    Efficient support for high-dimensional input: The integration of feature selection techniques (e.g., mRMR) reduces data dimensionality while preserving discriminative information, helping avoid vanishing kernel similarity and enabling practical deployment on NISQ devices.
  • $\bullet$
    Superior classification accuracy: Extensive experiments on multiple benchmark tasks (e.g., Credit Card, MNIST-5, FMNIST-4) and across various IBM quantum devices (e.g., Perth, Lagos, Torino) demonstrate that HaQGNN consistently outperforms several state-of-the-art quantum kernel baselines.
The remainder of this paper is organized as follows. Section II introduces background knowledge on quantum computing, kernel methods, and graph neural networks. Section III describes the proposed HaQGNN framework in detail. Section IV presents numerical results and comparative evaluations. Finally, Section V concludes the paper and discusses future research directions.
II Background
II.1 Quantum computation fundamental
A quantum bit (qubit) can exist in a linear superposition of the two computational basis states, $\left|0\right\rangle$ and $\left|1\right\rangle$, in contrast to a classical bit that can only be in a definite state of either 0 or 1. The state of a single qubit is generally represented as $\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle$, where $\alpha$ and $\beta$ are complex amplitudes satisfying the normalization condition $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$. For a system of multiple qubits, the overall quantum state is represented by the tensor product of individual qubit states, enabling the description of entangled states that cannot be factorized into separate 1-qubit components. Quantum gate, as a unitary matrix, transforms the quantum state to another and preserves the norm of the quantum state. Any unitary can be decomposed into a set of basis gates. The basis gate sets explored in this study are $S=\left\{R_{x},R_{z},CZ,I\right\}$ and $S^{\ast}=\left\{R_{z},X,CNOT,I\right\}$, which primarily consist of the 1-qubit gates $I$, $X$, $R_{x}$ and $R_{z}$, whose matrix representations are given as follows:
$I=\begin{pmatrix}1&&0\\ 0&&1\end{pmatrix}$ $X=\begin{pmatrix}0&&1\\ 1&&0\end{pmatrix}$ (1)
$R_{x}\left(\theta\right)=\begin{pmatrix}\cos\frac{\theta}{2}&-i\sin\frac{\theta}{2}\\ -i\sin\frac{\theta}{2}&\cos\frac{\theta}{2}\end{pmatrix}$ $R_{z}\left(\theta\right)=\begin{pmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{pmatrix}$ (2)
In addition, the 2-qubit gates $CNOT$ and $CZ$ are also included, with their corresponding matrix representations defined as:
$CNOT=\begin{pmatrix}1&&0&&0&&0\\ 0&&1&&0&&0\\ 0&&0&&0&&1\\ 0&&0&&1&&0\end{pmatrix}$ $CZ=\begin{pmatrix}1&&0&&0&0\\ 0&&1&&0&0\\ 0&&0&&1&0\\ 0&&0&&0&-1\end{pmatrix}$ (3)
II.2 Mechanism of quantum kernel
Kernel methods provide a powerful theoretical framework for performing nonlinear and nonparametric learning tasks, owing to their generality and strong interpretability [23]. Assume that both the training and test samples are drawn from the same input-output space $\mathcal{X}\times\mathcal{Y}$. The training dataset is denoted as $\mathcal{D}=\left\{\bm{x}_{i},y_{i}\right\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$, where $\bm{x}_{i}$ represents the input features of the $i$-th sample and $y_{i}$ denotes the corresponding label. The fundamental idea of traditional kernel methods is to map input data $\bm{x}_{i}$ from $\mathbb{R}^{d}$ to a higher-dimensional feature space $\mathbb{R}^{q}$, i.e., $\phi\left(\cdot\right):\mathbb{R}^{d}\to\mathbb{R}^{q}$ with $q\gg d$, such that data that are not linearly separable in the original space may become linearly separable in the transformed space. However, when $q$ is large, explicitly computing the mapped feature vectors $\phi\left(\bm{x}_{i}\right)$ becomes computationally expensive. To address this issue, kernel methods leverage the kernel trick, which enables implicit computation by constructing a kernel matrix $K\in\mathbb{R}^{n\times n}$, where each element is defined as the inner product of feature vectors in the high-dimensional space:
$\left[K\right]_{i,j}=\left[K\right]_{j,i}=\left\langle\phi\left(\bm{x}_{i}\right),\phi\left(\bm{x}_{j}\right)\right\rangle,\forall i,j\in\left[n\right]$ (4)
This approach enables efficient similarity computation with a time complexity of $\mathcal{O}\left(d\right)$, without explicitly performing the high-dimensional mapping. The aim of kernel learning is to infer a hypothesis function $h\left(\bm{x}_{i}\right)=\left\langle\bm{w}^{\ast},\phi\left(\bm{x}_{i}\right)\right\rangle$ with high test accuracy by utilizing the constructed kernel matrix $K$. The optimal parameter $\bm{w}^{\ast}$ is obtained by minimizing the following loss function:
$\mathcal{L}=\lambda\left\langle\bm{w},\bm{w}\right\rangle+\sum_{i=1}^{n}\left(\left\langle\bm{w},\phi\left(\bm{x}_{i}\right)\right\rangle\right)^{2}$ (5)
The performance of kernel methods heavily depends on the choice of the feature mapping function $\phi\left(\cdot\right)$, or equivalently, the kernel function $\kappa\left(\cdot,\cdot\right)=\left\langle\phi\left(\cdot\right),\phi\left(\cdot\right)\right\rangle$. Various kernel functions have been proposed to accommodate different task scenarios, such as the radial basis function (RBF) kernel, Gaussian kernel, circular kernel, and quantum kernel.
Compared to classical kernel methods, the primary distinction of quantum kernel approaches lies in the implementation of the feature mapping function $\phi\left(\cdot\right)$. For an $N$-qubit quantum kernel model, the feature mapping of a sample $\bm{x}_{i}$ is realized through a parameterized quantum circuit $U_{c}\left(\bm{x}_{i},\bm{\theta}\right)$ applied to the initial quantum state $\left|0\right\rangle^{\otimes N}$, resulting in the quantum state $\left|\varphi\left(\bm{x}_{i},\bm{\theta}\right)\right\rangle=U_{c}\left(\bm{x}_{i},\bm{\theta}\right)\left|0\right\rangle^{\otimes N}$. Here, $\bm{\theta}$ denotes trainable parameters and the corresponding density matrix is denoted as $\rho\left(\bm{x}_{i},\bm{\theta}\right)=\left|\varphi\left(\bm{x}_{i},\bm{\theta}\right)\right\rangle\left\langle\varphi\left(\bm{x}_{i},\bm{\theta}\right)\right|$. The objective of quantum kernel learning is to utilize the quantum kernel matrix $W\left(\bm{x},\bm{\theta}\right)$, defined as
$\left[W\left(\bm{x},\bm{\theta}\right)\right]_{i,j}=Tr\left(\rho\left(\bm{x}_{i},\bm{\theta}\right)\rho\left(\bm{x}_{j},\bm{\theta}\right)\right),\forall i,j\in\left[n\right]$ (6)
to infer a hypothesis as formulated in Eq. 5, with the goal of achieving higher test accuracy compared to classical kernel methods. Therefore, the performance of a quantum kernel largely depends on the design of the parameterized circuit $U_{c}\left(\cdot,\bm{\theta}\right)$. It is worth noting that the implementation of $U_{c}\left(\cdot,\bm{\theta}\right)$ is flexible and can take various forms, such as circuits constructed with predefined architectures, hardware efficient circuits, or circuits randomly generated from a specified quantum gate set.
II.3 Graph neural networks
Graph Neural Networks (GNNs) have become a powerful framework for learning over graph-structured data, where relationships among entities play a critical role in the underlying data distribution [32]. By iteratively aggregating and transforming information from neighboring nodes through message-passing mechanisms, GNNs are capable of capturing both local topology and global structural dependencies [33]. This capacity to model complex relational patterns has led to their widespread adoption across a variety of domains, including social network analysis [34], molecular property prediction [35, 36, 37], and recommender systems [38, 39]. Notably, variants such as Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and GraphSAGE have significantly advanced the state-of-the-art in graph representation learning tasks [40, 41, 42, 43, 44]. In the context of quantum computing, quantum circuits can be naturally represented as graphs, where qubits correspond to nodes and quantum gates or couplings define the edges. This structural analogy provides a compelling motivation for applying GNNs to quantum systems, enabling efficient learning and prediction tasks grounded in the intrinsic connectivity of quantum architectures.
III Method
In this section, we present the technical details of our proposed hardware-aware quantum kernel design framework. An overview of the framework is illustrated in Fig. 1, which consists of four main steps:
Step 1: Based on the hardware topology and noise characteristics of the quantum device, a set of qubits is selected such that they exhibit favorable connectivity and minimal susceptibility to noise. A large number of candidate quantum circuits are then generated on the selected qubits, following the supported native gate set and qubit connectivity of the target device.
Step 2: A graph neural network (GNNs-1) is employed to predict the ”Probability of Successful Trials (PST)”, a metric correlated with circuit fidelity. This step enables the early rejection of low-fidelity candidate circuits, thereby filtering out circuits with potentially higher fidelity.
Step 3: A second graph neural network (GNNs-2) is designed to estimate the kernel-target alignment (KTA), a surrogate metric for classification accuracy, over the high-fidelity circuits identified in Step 2. The top-$k$ circuits with the highest predicted KTA values are selected for further refinement.
Step 4: A classical optimizer is employed to train the parameters of the top-$k$ high-KTA circuits with the high PST values from scratch, and the circuit with the best task-specific classification accuracy is selected as the final quantum kernel circuit.
We divide the hardware-aware quantum kernel design into three parts: candidate circuit generation (Step 1, Section III.1), metric prediction based on graph neural networks (Step 2-3, Section III.2), circuit training and selection (Step 4, Section III.3).
III.1 Candidate circuit generation
The process of generating candidate quantum circuits is primarily divided into three stages: 1) A specific number of qubits with preserved topological connectivity are selected from the quantum device. This selection process takes into account the noise characteristics of individual qubits and their couplings, aiming to identify a subgraph of the device topology with the lowest average noise. Such the selection strategy improves hardware efficiency and enhances the fidelity of the generated circuits. 2) Candidate circuits are generated based on the selected qubit and the native gate set supported by the quantum device. During the circuit generation process, each inserted quantum gate must act on qubits that are physically connected on the hardware. This ensures that qubit mapping is implicitly achieved during circuit construction, thereby avoiding the insertion of costly SWAP gates during the compilation phase. As a result, the overall circuit depth is reduced, and noise accumulation is mitigated. 3) In the candidate quantum circuits, parameterized quantum gates are selected at different positions to enable flexible data embedding, which helps improve the performance of the quantum kernel.
III.1.1 Subgraph selection in device topology
Due to the non-negligible differences in noise levels across individual qubits and qubit couplings within quantum devices, we perform a selection process to identify a set of $N$ qubits that are less affected by noise. This selection aims to improve hardware efficiency and maintain higher circuit fidelity. Fig. 2 illustrates examples of selected subgraphs with $N=4,5,7$ and $8$ extracted from the 133-qubit IBM Torino topology.
Figure 2: Subgraphs with $N=4,5,7$ and $8$ selected from the 133-qubit IBM Torino topology.
Selecting a subgraph with the desired $N$ qubits requires a comprehensive consideration of several hardware-specific factors in the quantum device topology, including 1-qubit gate errors, readout errors, 2-qubit gate errors between coupled qubits, and the degree of connectivity among qubits. The target is to identify a subgraph with high gate fidelity and high qubit connectivity, while maintaining low readout errors. By incorporating device-specific noise information, the subgraph selection strategy ensures that the subsequently generated quantum circuits are noise-aware, with detailed procedures provided in Appendix A.
III.1.2 Generating candidate circuits based on the subgraph
After selecting a high-fidelity subgraph, denoted as $G$, we generate $M$ candidate quantum circuits based on the structure of $G$ and the set of quantum gates supported by the device, denoted as $S$. This process ensures that the generated circuits are hardware-aware. For example, as shown in Step 1 of Fig. 1, candidate circuits are generated from the subgraph of the IBM Torino using the gate set $S=\left\{R_{x},R_{z},CZ,I\right\}$. The detailed procedure for candidate circuit generation is provided in Algorithm III.1.2. During the candidate circuit generation process, we guide the selection of 2-qubit gate between different coupled qubits based on their gate errors. Specifically, the probability of inserting a 2-qubit gate between qubits $q_{i}$ and $q_{j}$, denoted as $P\left(q_{i},q_{j}\mid e_{j}\right)$, is defined as follows:
$\begin{array}[]{c}P\left(q_{i},q_{j}\mid e_{j}\right)=\frac{1/e_{j}}{{\textstyle\sum_{j}1/e_{j}}}\end{array}$ (7)
where the 2-qubit gate error $e_{j}$ corresponds to the edge $q_{i}q_{j}$ in subgraph $G$. 1-qubit gate errors are not considered because 2-qubit gate errors are typically an order of magnitude higher and have a more significant impact on circuit performance. Moreover, since the construction of candidate circuits is guided by noise information, the resulting circuits are inherently noise-aware. By repeatedly executing the procedure in Algorithm III.1.2, a large number of hardware- and noise-aware candidate quantum circuits can be generated, laying the foundation for subsequent GNN-based training and the identification of optimal quantum circuits.
Algorithm 1: Generate candidate circuit
1:Device subgraph $G\left(\left\{V_{i}\right\},\left\{E_{j}\right\}\right)$; 2-qubit gate error $\left\{e_{j}\right\}$; The gate set $S$; The number of gates $P$.
2:Candidate circuit $C$.
3:for $k$ in range($P$do
4:     Sample the gate $g_{k}$ from $S$ on average $\triangleright$ Sample a gate to be added, e.g., $R_{x}$, $R_{z}$, or $CZ$.
5:     if $g_{k}$ is a 1-qubit gate then
6:         Select a qubit $q_{i}\in\left\{V_{i}\right\}$ with average probability from subgraph $G$triangleright Select a qubit from the subgraph as the target qubit of the quantum gate.

### v2

@@
-Hardware-Aware Quantum Kernel Design Based on Graph Neural Networks
+HaQGNN: Hardware-Aware Quantum Kernel Design Based on Graph Neural Networks
 Abstract
-Quantum kernel methods have emerged as a promising direction in quantum machine learning (QML), offering a principled way to map classical data into high-dimensional quantum Hilbert spaces.
-While conceptually powerful, designing effective quantum kernels that adapt to both the target task and the constraints of near-term quantum hardware remains a nontrivial challenge.
-In this work, we propose HaQGNN, a hardware-aware quantum kernel design framework that integrates quantum device topology, noise characteristics, and graph neural networks (GNNs) to evaluate and select task-relevant quantum circuits.
-By predicting surrogate metrics related to fidelity and kernel performance, HaQGNN enables efficient circuit screening at scale.
-Feature selection is further incorporated to improve compatibility with limited-qubit systems and mitigate kernel degradation.
-Extensive experiments on three benchmark datasets, Credit Card, MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms existing quantum kernel baselines in terms of classification accuracy.
-Our work highlights the potential of learning-based and hardware-aware strategies for advancing practical quantum kernel design on NISQ devices.
+Designing effective quantum kernels is a central challenge in Quantum Machine Learning (QML), particularly under the limitations of Noisy Intermediate-Scale Quantum (NISQ) devices with a limited number of qubits, error-prone gate operations, and restricted qubit connectivity.
+To address this, we propose HaQGNN, a hardware-aware quantum kernel design method that integrates quantum device topology, noise characteristics, and Graph Neural Networks (GNNs) to evaluate and select task-relevant quantum circuits that define quantum kernels.
+First, each quantum circuit is represented as a directed acyclic graph that encodes hardware-specific features, including gate types, target qubits, and noise characteristics.
+Next, two GNNs are trained to predict surrogate metrics, Probability of Successful Trials (PST) and Kernel-Target Alignment (KTA), for fast and accurate fidelity and performance estimation.
+Additionally, feature selection is further incorporated to reduce input dimensionality and improve compatibility with limited-qubit devices.
+Finally, extensive experiments on three benchmark datasets, Credit Card (CC), MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms existing baselines in terms of classification accuracy.
+Our results highlight the potential of learning-based and hardware-aware strategies for advancing practical quantum kernel design on near-term quantum hardware.
 preprint: APS/123-QED
@@
 Quantum computing presents a compelling prospect for revolutionizing machine learning tasks by exploiting the fundamental properties of quantum mechanics, such as superposition, entanglement, and high-dimensional Hilbert spaces [1, 2].
-Quantum machine learning (QML), as a important class of quantum algorithms tailored for the Noisy Intermediate-Scale Quantum (NISQ) era [3], has demonstrated practical applicability in diverse real-world tasks, including classification [4, 5, 6, 7, 8], generative modeling [9, 10, 11, 12], and learning physical systems [13, 14].
+Quantum Machine Learning (QML), as an important class of quantum algorithms tailored for the Noisy Intermediate-Scale Quantum (NISQ) era [3], has demonstrated practical applicability in diverse real-world tasks, including classification [4, 5, 6, 7, 8], generative modeling [9, 10, 11, 12], and learning physical systems [13, 14].
 Among various QML approaches, quantum kernel methods have attracted increasing attention due to their solid theoretical foundation and interpretability [15, 16, 17].
 By embedding classical data into quantum states through parameterized quantum circuits, quantum kernels define implicit feature maps capable of capturing complex nonlinear relationships beyond classical means.
-Despite recent demonstrations of quantum kernel learning on real quantum devices, realizing practical quantum advantage remains challenging, especially under the constraints of Noisy Intermediate-Scale Quantum (NISQ) devices [18, 19, 20].
+Despite recent demonstrations of quantum kernel learning on real quantum devices, realizing practical quantum advantage remains challenging, especially under the constraints of NISQ devices [18, 19, 20].
 In particular, unsuitable quantum circuit designs may lead to vanishing kernel similarities [21], reduced expressivity, and degraded generalization [22, 23], severely limiting the performance of quantum kernels.
@@
 However, most existing approaches either ignore hardware constraints such as qubit connectivity and gate noise, or suffer from excessive computational costs in searching large quantum circuit spaces.
-Figure 1: An overview of handware-aware quantum kernel design.
-To address these challenges, we propose a hardware-aware quantum kernel design framework based on Graph Neural Networks (HaQGNN), which efficiently integrates quantum hardware topology and device noise characteristics into the design of quantum feature maps, as illustrated in Fig.
+To address these challenges, we propose HaQGNN, a hardware-aware quantum kernel design based on Graph Neural Networks, which efficiently integrates quantum hardware topology and device noise characteristics into the design of quantum feature maps, as illustrated in Fig.
 1.
@@
 Rather than formulating quantum kernel design as a joint optimization over circuit structure and parameters, our approach decouples the process by first generating hardware-aware candidate circuits and then employing graph neural networks to evaluate their potential performance.
-Leveraging the graph structure of quantum circuits, we employ GNNs to predict two surrogate metrics: the ”Probability of Successful Trials (PST)” [31], which reflects circuit fidelity under hardware noise, and kernel-target alignment (KTA), which correlates with downstream classification accuracy.
-The HaQGNN framework consists of four steps: (1) candidate circuit generation based on hardware topology and native gate sets; (2) PST prediction via GNNs-1 to filter out low-fidelity circuits; (3) KTA prediction via GNNs-2 to select high-performing kernels; and (4) parameter optimization of the top-$k$ circuits for final task-specific performance.
+Leveraging the graph structure of quantum circuits, we employ GNNs to predict two surrogate metrics: the Probability of Successful Trials (PST) [31], which reflects circuit fidelity under hardware noise, and Kernel-Target Alignment (KTA), which correlates with downstream classification accuracy.
+The HaQGNN method consists of four steps: (1) candidate circuit generation based on hardware topology and native gate sets; (2) PST prediction via GNNs-1 to filter out low-fidelity circuits; (3) KTA prediction via GNNs-2 to select high-performing kernels; and (4) parameter optimization of the top-$k$ circuits for final kernel selection.
 Furthermore, we incorporate feature selection to reduce input dimensionality, mitigating vanishing similarity issues and enhancing compatibility with NISQ devices.
-To the best of our knowledge, this is the first work to propose a GNN-based, hardware-aware framework that evaluates and selects quantum kernel circuits by jointly considering hardware topology, noise characteristics, and circuit structure, followed by task-specific parameter optimization.
+To the best of our knowledge, this is the first work to propose a GNN-based, hardware-aware method that evaluates and selects quantum kernel circuits by jointly considering hardware topology, noise characteristics, and circuit structure, followed by task-specific parameter optimization.
 Extensive simulations on IBM quantum devices demonstrate the superiority of HaQGNN over several state-of-the-art quantum kernel baselines across multiple classification tasks and device configurations.
-Our proposed HaQGNN framework offers several distinctive advantages that collectively enable efficient quantum kernel design for NISQ devices:
+Our proposed HaQGNN method offers several distinctive advantages that collectively enable efficient quantum kernel design for NISQ devices:
   • $\bullet$
-    Hardware-awareness: HaQGNN explicitly incorporates real-device topology and noise information during both circuit generation and performance prediction, enhancing compatibility and circuit fidelity across different quantum hardware backends.
+    Hardware-awareness: HaQGNN explicitly incorporates real-device topology and noise information during both circuit generation and performance prediction, enhancing circuit performance across different quantum hardware backends.
   • $\bullet$
-    Two-stage GNN-based performance estimation: By employing two lightweight graph neural networks (GNNs-1 and GNNs-2), the framework accurately predicts fidelity (PST) and classification potential (KTA) of circuits with minimal simulation cost, achieving speedups of over 380× and 58000× compared to traditional metrics.
+    Dual-GNN-based performance prediction: By employing two graph neural networks (GNNs-1 and GNNs-2), the HaQGNN accurately predicts fidelity (PST) and classification accuracy (KTA) of circuits with low simulation cost, achieving speedups of over 380× and 58000× compared to direct calculation.
   • $\bullet$
-    Noise-robust candidate selection: By filtering out low-fidelity circuits early in the pipeline, HaQGNN effectively mitigates the impact of gate errors and decoherence, leading to more reliable kernel performance on noisy devices.
+    Noise-aware candidate filtering: By filtering out low-fidelity circuits early in the pipeline, HaQGNN effectively mitigates the impact of gate errors and decoherence, leading to more reliable kernel performance on noisy devices.
   • $\bullet$
@@
   • $\bullet$
-    Superior classification accuracy: Extensive experiments on multiple benchmark tasks (e.g., Credit Card, MNIST-5, FMNIST-4) and across various IBM quantum devices (e.g., Perth, Lagos, Torino) demonstrate that HaQGNN consistently outperforms several state-of-the-art quantum kernel baselines.
+    Superior classification accuracy: Extensive experiments on multiple benchmark tasks (e.g., Credit Card (CC), MNIST-5, FMNIST-4) and across various IBM quantum devices (e.g., Perth, Lagos, Torino) demonstrate that HaQGNN consistently outperforms several state-of-the-art baselines.
 The remainder of this paper is organized as follows.
 Section II introduces background knowledge on quantum computing, kernel methods, and graph neural networks.
-Section III describes the proposed HaQGNN framework in detail.
+Section III describes the proposed HaQGNN method in detail.
 Section IV presents numerical results and comparative evaluations.
@@
 By iteratively aggregating and transforming information from neighboring nodes through message-passing mechanisms, GNNs are capable of capturing both local topology and global structural dependencies [33].
-This capacity to model complex relational patterns has led to their widespread adoption across a variety of domains, including social network analysis [34], molecular property prediction [35, 36, 37], and recommender systems [38, 39].
+This capacity to model complex relational patterns has led to their widespread adoption across a variety of domains, including social network analysis [34], molecular property prediction [35, 36, 37, 38], and recommender systems [39, 40].
 Notably, variants such as Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and GraphSAGE have significantly advanced the state-of-the-art in graph representation learning tasks [41, 42, 43, 44, 45].
@@
 III Method
-In this section, we present the technical details of our proposed hardware-aware quantum kernel design framework.
-An overview of the framework is illustrated in Fig.
-1, which consists of four main steps:
-Step 1: Based on the hardware topology and noise characteristics of the quantum device, a set of qubits is selected such that they exhibit favorable connectivity and minimal susceptibility to noise.
+Figure 1: An overview of HaQGNN.
+In this section, we present the technical details of our proposed HaQGNN. An overview of this method is illustrated in Fig. 1, which consists of four main steps:
+Step 1 Candidate Circuit Generation: Based on the hardware topology and noise characteristics of the quantum device, a set of qubits is selected such that they exhibit favorable connectivity and low susceptibility to noise.
 A large number of candidate quantum circuits are then generated on the selected qubits, following the supported native gate set and qubit connectivity of the target device.
-Step 2: A graph neural network (GNNs-1) is employed to predict the ”Probability of Successful Trials (PST)”, a metric correlated with circuit fidelity.
+Step 2 PST Prediction: A graph neural network (GNNs-1) is employed to predict the PST, a metric correlated with circuit fidelity.
 This step enables the early rejection of low-fidelity candidate circuits, thereby filtering out circuits with potentially higher fidelity.
-Step 3: A second graph neural network (GNNs-2) is designed to estimate the kernel-target alignment (KTA), a surrogate metric for classification accuracy, over the high-fidelity circuits identified in Step 2.
+Step 3 KTA Prediction: A second graph neural network (GNNs-2) is designed to estimate the KTA, a surrogate metric for classification accuracy, over the high-fidelity circuits identified in Step 2.
 The top-$k$ circuits with the highest predicted KTA values are selected for further refinement.
-Step 4: A classical optimizer is employed to train the parameters of the top-$k$ high-KTA circuits with the high PST values from scratch, and the circuit with the best task-specific classification accuracy is selected as the final quantum kernel circuit.
-We divide the hardware-aware quantum kernel design into three parts: candidate circuit generation (Step 1, Section III.1), metric prediction based on graph neural networks (Step 2-3, Section III.2), circuit training and selection (Step 4, Section III.3).
+Step 4 Final Training and Kernel Selection: A classical optimizer is employed to train the parameters of the top-$k$ high-KTA circuits with the high PST values from scratch, and the circuit with the best task-specific classification accuracy is selected as the final quantum kernel.
+To facilitate presentation, we organize the four-step HaQGNN method into three parts: candidate circuit generation (Step 1, Section III.1), metrics prediction based on graph neural networks (Step 2-3, Section III.2), final training and kernel selection (Step 4, Section III.3).
 III.1 Candidate circuit generation
@@
 This selection process takes into account the noise characteristics of individual qubits and their couplings, aiming to identify a subgraph of the device topology with the lowest average noise.
-Such the selection strategy improves hardware efficiency and enhances the fidelity of the generated circuits.
+Such a selection strategy improves hardware efficiency and enhances the fidelity of the generated circuits.
 2) Candidate circuits are generated based on the selected qubit and the native gate set supported by the quantum device.
@@
 6:         Select a qubit $q_{i}\in\left\{V_{i}\right\}$ with average probability from subgraph $G$triangleright Select a qubit from the subgraph as the target qubit of the quantum gate.
+7:         Append $g_{k}\left(q_{i}\right)$ to $C$
+8:     else if $g_{k}$ is a 2-qubit gate then
+9:         Select a pair of qubits $q_{i},q_{j}$ with the probability $P\left(q_{i},q_{j}\mid e_{j}\right)$ and the corresponding edge $q_{i}q_{j}\in\left\{E_{j}\right\}$ from subgraph $G$triangleright 2-qubit gate placement is guided by the associated error $e_{j}$.
+10:         Append $g_{k}\left(q_{i}q_{j}\right)$ to $C$
+11:     end if
+12:end for
+13:return $C$
+III.1.3 Data embedding in circuits
+Figure 3: The randomized data embedding strategy.
+Inspired by the advantages of randomized data embedding over fixed embedding discussed in [7], we adopt a simple strategy for data embedding, as illustrated in Fig.
+3: in each quantum circuit, a subset of parameterized 1-qubit gates is randomly selected to embed the data, with each selected gate corresponding to one dimension of the input data.
+To ensure that the entire input can be encoded, the number of embedding gates should be greater than or equal to the data dimensionality.
+However, for image classification tasks, the data dimensionality is typically huge; for instance, each sample in the MNIST dataset has 784 dimensions.
+Embedding such high-dimensional data requires a large number of quantum gates, leading to circuits with a large number of qubits and high circuit depths.
+Due to the vanishing similarity issue, larger qubit numbers and deeper circuits degrade the performance of candidate circuits.
+To address this, we apply a feature selection method in the preprocessing stage to reduce input dimensionality.
+Specifically, we use the Max-Relevance and Min-Redundancy (mRMR) algorithm to reduce the data dimensionality from $d$ to $p$, where $p\ll d$.
+This significantly decreases the number of gates $P$ required for candidate circuit construction.
+This approach not only mitigates the vanishing similarity issue but also enables the effective utilization of high-dimensional data on NISQ devices.
+III.2 Metrics prediction based on graph neural networks
+To predict the PST of candidate circuits under device noise and the corresponding performance metric KTA, we design two separate GNNs, namely GNNs-1 and GNNs-2.
+These two networks share an overall similar architecture but differ in certain details due to the differences in their prediction targets.
+In the following, we provide a detailed description of the GNNs construction process, covering the aspects of construction of the dataset, graph construction, node features, target values, and construction of GNNs.
+III.2.1 Construction of the dataset
+The primary motivation for constructing GNNs predictors is to enable rapid predictions on a large set of candidate circuits after training with only a small amount of data.
+To build the dataset for training the GNNs, we adopt a random sampling strategy: a subset of circuits is randomly selected from the previously generated large pool of candidate circuits.
+To ensure the completeness of each circuit, we initialize all qubits in the state $\left|0\right\rangle$ and include measurement operations within the circuits, as illustrated in the left panel of Fig.
+4.
+This setup also lays the foundation for incorporating device-specific noise characteristics, such as readout errors, into the node features.
+III.2.2 Graph construction
+Figure 4: The conversion of quantum circuit to graph structure.
+Motivated by the intuitive similarity between quantum circuits and graph structures, we represent quantum circuits using directed acyclic graphs (DAGs).
+In this representation, each node corresponds to a qubit, a quantum gate, or a measurement operation, while the edges capture the temporal sequence of operations, progressing from left to right.
+As shown in Fig.
+4, we perform the transformation from circuits generated on IBM Torino into graph structures, with particular care taken to represent each 2-qubit gate as a single graph node.
+III.2.3 Node features
+For each node in the graph structure, we construct a feature vector to represent its associated properties.
+Features include index, node type, target qubit, tag, relaxation time ($T_{1}$) and dephasing time ($T_{2}$) of the target qubits, gate error, and readout error.
+The specific node features for GNNs-1 and GNNs-2 are shown in Fig.
+5.
+Figure 5: Node feature vectors of GNNs-1 and GNNs-2.
+The feature vectors for both GNNs-1 and GNNs-2 have a length of 31.
+The first part of the feature vector encodes the index, which is used for node sorting.
+The next three parts are one-hot encodings representing the node type, target qubit, and tag, respectively.
+The node type distinguishes different types of nodes, including initial input qubit, measurement, $R_{x}$ gate, $R_{z}$ gate, $CZ$ gate, and identity ($I$) gate.
+For example, the one-hot code $000010$ indicates that the node corresponds to a $CZ$ gate.
+The target qubit labels the index of the qubit affected by the operation, and its length can be expanded according to practical needs.
+The tag distinguishes whether a gate is used for data embedding, switching from 0 to 1 for embedding gates.
+In addition to capturing the topological connections between circuits, another advantage of GNNs is their ability to incorporate hardware noise information, which is primarily achieved through node feature vectors.
+In our method, the last seven elements of each node’s feature vector are dedicated to describing the noise characteristics of the quantum hardware.
+During model training, GNNs-1 incorporates these noise features to predict fidelity-related metric and enable early rejection of poor-performing circuits.
+In contrast, GNNs-2 focuses on predicting performance-related metric, with the last 7 elements of its feature vectors zero-filled to exclude noise information.
+The noise parameters are based on real quantum devices such as IBM Torino, IBM Perth, and IBM Nairobi.
+Specifically, the first 4 elements represent the relaxation time ($T_{1}$) and dephasing time ($T_{2}$) for the two involved qubits, the 5th element encodes the gate operation error, and the final two elements capture the readout errors.
+For nodes without associated noise characteristics, the corresponding values are set to zero.
+For instance, measurement nodes have their $T_{1}$, $T_{2}$, and gate error fields set to zero.
+Fig.
+5 shows the complete feature vector representation of a $CZ$ gate node for both GNNs-1 and GNNs-2.
+III.2.4 Target values
+To evaluate the impact of noise on different quantum circuits, the concept of fidelity is introduced, which quantifies the closeness between two quantum states.
+In noisy quantum computations, fidelity is typically used to measure the difference between the quantum state generated by a noisy circuit and that generated by a noiseless classical simulation.
+A significant deviation indicates that the circuit is severely affected by noise, and its output is likely to be inaccurate.
+However, directly obtaining the fidelity of a quantum circuit requires complex quantum state tomography, which is computationally costly.
+To address this issue, we adopt the ”Probability of Successful Trials (PST)” proposed in [31] as a proxy for fidelity, which has been shown to provide accurate fidelity estimates.
+The PST is specifically defined as follows:
+$PST=\frac{T_{initial}}{T_{total}}$ (8)
+where $T_{initial}$ denotes the number of trials with an output identical to the initial state $\left|0\right\rangle^{\otimes N}$, and $T_{total}$ represents the total number of trials.
+Instead of directly computing fidelity, we concatenate each circuit with its inverse and apply the combined circuit to the initial state $\left|0\right\rangle^{\otimes N}$.
+The proportion of measurements yielding the initial state is then recorded to obtain the PST, as illustrated in Fig.
+6.
+The computed PST is used as the target label for training the GNNs-1 model.
+Figure 6: The computation process of PST.
+GNNs-2 is designed to predict a metric known as Kernel-Target Alignment (KTA), which has been shown to be positively correlated with training accuracy.
+The relationship between KTA and training accuracy is analyzed in [23], where KTA is validated as a reliable surrogate for training accuracy.
+Before introducing KTA, we first briefly review the concept of kernel alignment, which was first introduced by [46].
+Kernel alignment is a similarity measure between two kernel matrices $K_{1}$ and $K_{2}$:
+$KA\left(K_{1},K_{2}\right)=\frac{\left\langle{K_{1},K_{2}}\right\rangle_{F}}{\sqrt{\left\langle{K_{1},K_{1}}\right\rangle_{F}\left\langle{K_{2},K_{2}}\right\rangle_{F}}}$ (9)
+where $\left\langle K_{1},K_{2}\right\rangle_{F}$ denotes the Frobenius inner product between two matrices, which is given by
+$\left\langle K_{1},K_{2}\right\rangle_{F}=\sum_{i=1}^{i=l}\sum_{j=1}^{j=l}k_{1}\left(x_{i},x_{j}\right)k_{2}\left(x_{i},x_{j}\right)$ (10)
+where $l$ denotes the order of the matrix.
+For classification purposes, we can define an ideal target matrix as $K^{\ast}=yy^{T}$, where $y=\left(y_{1},y_{2},\cdots,y_{l}\right)^{T}$ is the vector of labels for the training set $D$.
+The kernel-target alignment is then defined as the kernel alignment of the kernel matrix $K$, which can be generated by the quantum kernel, and $K^{\ast}$:
+$KTA\left(K\right)=KA\left(K,K^{\ast}\right)=\frac{\left\langle K,yy^{T}\right\rangle_{F}}{\sqrt{\left\langle K,K\right\rangle_{F}\left\langle yy^{T},yy^{T}\right\rangle_{F}}}$ (11)
+Since $\left\langle A,B\right\rangle=Tr\left(A^{T}B\right)$ and $K^{T}=K$, substituting them into Eq. 11 yields the following formula:
+$KTA\left(K\right)=\frac{Tr\left(Kyy^{T}\right)}{\sqrt{Tr\left(K^{2}\right)Tr\left(\left(yy^{T}\right)^{2}\right)}}=\frac{y^{T}Ky}{l\sqrt{Tr\left(K^{2}\right)}}$ (12)
+Here, $l$ also denotes the number of elements in the label vector $y$.
+For multi-class classification problems (including binary classification), the ideal target matrix $K^{\ast}$ can be constructed using the following formula [47]:
+$\left[K^{\ast}\right]_{i,j}=\begin{cases}1&y_{i}=y_{j}\\ \frac{-1}{c-1}&y_{i}\neq y_{j}\end{cases}$ (13)
+where $c$ ($c\geq 2$) is the number of classes.
+Given that KTA is a reliable surrogate for classification accuracy and can be efficiently computed, we adopt it as a performance metric for quantum circuits and use it as the target label for training GNNs-2.
+Figure 7: The architectures of the GNNs-1 and GNNs-2.
+III.2.5 Construction of GNNs
+We design similar graph neural network (GNN) architectures for GNNs-1 and GNNs-2, as illustrated in Fig.
+7.
+Both models adopt a four-layer structure, where each layer is followed by an attention mechanism.
+The incorporation of attention allows the network to dynamically assign weights to edges during message passing, thereby enabling more effective capture of circuit characteristics.
+Between layers, LeakyReLU is used as the activation function.
+After graph-based feature extraction, we perform average pooling over the features of key nodes (excluding input and measurement nodes).
+The resulting pooled vector is then concatenated with the output of a three-layer fully connected network that processes the circuit’s global feature vector.The concatenated representation is subsequently passed through a multilayer perceptron (MLP) to generate the final prediction for either GNNs-1 or GNNs-2, as shown on the right side of Fig.
+7.
+While GNNs-1 and GNNs-2 share the same overall architecture, they differ in the composition of their global feature vector.
+Specifically, GNNs-1 takes as global features $g_{1}$: the count of gate, $k_{1}$: the count of 2-qubit gate, and $z_{1}$: depth of the circuit.
+In contrast, GNNs-2 uses $g_{2}$: the count of gate, $k_{2}$: the count of 2-qubit gate, and $z_{2}$: the count of embedding gate.
+These global feature vectors are selected based on the specific prediction targets of each model, allowing for improved prediction performance.
+III.3 Final training and kernel selection
+This step aims to identify the most effective quantum kernel for the target classification task.
+Based on the results of Step 2 and Step 3, we select the top-$k$ quantum circuits that jointly exhibit high predicted values of PST and KTA.
+These $k$ circuits are then applied to the target classification task for full training.
+For each selected circuit, the trainable parameters are optimized from scratch using the classical Adam optimizer, while keeping the circuit structure fixed.
+After training, we evaluate the classification accuracy of each circuit on the test set and select the one with the highest accuracy as the final quantum kernel.
+IV Evaluation
+In this section, we conduct a series of experiments to thoroughly evaluate the performance of HaQGNN.
+The experimental evaluation is divided into two main parts.
+The first part assesses the predictive performance and computational efficiency of the trained graph neural networks, GNNs-1 and GNNs-2 (Section IV.1).
+The second part demonstrates the performance advantages of HaQGNN over recent state-of-the-art methods across different datasets through simulation studies (Section IV.2).
+IV.1 Predictive performance of the graph neural networks
+Model and Training Setups.
+In the default setup, both GNNs-1 and GNNs-2 adopt a 4-layer Graph Neural Network architecture with an attention mechanism integrated into each layer.
+This allows the model to dynamically assign weights to neighboring nodes during message passing, following the design of Graph Attention Networks (GATs).
+At the final GNN layer, both GNNs-1 and GNNs-2 perform average pooling over key nodes (excluding input and measurement nodes) to obtain a 31-dimensional pooled vector.
+This pooled vector is then concatenated with a global feature vector obtained from a three-layer fully connected network, where each hidden layer has a dimensionality of 12.
+The resulting concatenated vector is subsequently passed through a multilayer perceptron (MLP) with hidden layers of size 256, 128, and 64 to produce the final prediction.
+The activation function used in our model is LeakyReLU with a hyperparameter $\alpha=0.02$.
+Training is performed using the Adam optimizer with a learning rate of 0.01, a batch size of 512, and the mean squared error (MSE) as the loss function.
+The network is trained for 200 epochs.
+During data preprocessing, we apply Min-Max normalization to the node features across the entire dataset.
+The predictive performance is evaluated using the coefficient of determination ($R^{2}$), defined as:
+$R^{2}=1-\frac{{\textstyle\sum_{i=1}^{N}}\left(y_{i}-\hat{y}_{i}\right)^{2}}{{\textstyle\sum_{k=1}^{N}}\left(y_{i}-\bar{y}\right)^{2}}$ (14)
+where $\hat{y}_{i}$ is the output prediction associated with the ground truth target value $y_{i}$, $\bar{y}$ is the average of the target values, $N$ is the number of circuits in the test set.
+Dataset Setup.
+Following the methodology described in Section III.1, we generate a total of 50000 candidate circuits and randomly select 5000 of them for training GNNs-1 and GNNs-2.
+Specifically, 4000 circuits are used for training and the remaining 1000 for testing.
+Although we adopt surrogate metrics (PST and KTA) to significantly reduce the cost of label generation, computing these metrics on a large scale remains time-consuming.
+This trade-off is further illustrated in our subsequent runtime comparisons.
+For GNNs-1, the noise models used in training are derived from real quantum hardware, including several 7-qubit devices (IBM Perth, IBM Lagos, IBM Nairobi, and IBM Jakarta) and the 133-qubit IBM Torino.
+The considered noise types include gate errors, relaxation time ($T_{1}$), dephasing time ($T_{2}$), and readout errors.
+In addition, the gate set of all 7-qubit quantum devices is $S^{\ast}=\left\{R_{z},X,CNOT,I\right\}$.
+To evaluate the predictive performance of GNNs-1 and GNNs-2 in PST and KTA prediction, respectively, we conduct simulations under varying numbers of qubits: $N=4,5,7$ and $8$.
+As shown in Fig.
+8, the simulations are performed on quantum devices corresponding to these configurations: IBM Perth for $N=4$, and IBM Torino for $N=5,7$ and $8$.
+In each case, 100 samples are randomly selected from the test set for visualization, while the coefficient of determination ($R^{2}$) is computed using the entire test set.
+(a) Scatter plot of predicted PST against ground truth values
+(b) Scatter plot of predicted KTA against ground truth values
+Figure 8: Scatter plots of predicted PST and KTA values with $N=4,5,7$ and $8$.
+The results in Fig.
+8 demonstrate that both GNNs-1 and GNNs-2 achieve high prediction accuracy, with $R^{2}$ exceeding 0.97 for PST and 0.95 for KTA.
+The high accuracy of GNNs-1 in predicting PST under hardware noise is attributed to the inclusion of noise information in the node feature vectors, which enhances the model’s ability to learn noise-aware representations.
+Moreover, we observe that KTA values are tightly clustered, while PST values exhibit a broader distribution.
+This implies that quantum kernels with similar KTA values, i.e., theoretically similar classification performance, can exhibit significantly different results on real hardware due to noise.
+This observation further justifies our strategy of early rejection of low-fidelity candidate circuits based on PST prediction.
+Figure 9: Comparison of runtime of prediction and direct calculation for PST and KTA.
+To further highlight the efficiency of our approach, we evaluate the computational speedup achieved by using GNNs-1 and GNNs-2 for circuit performance estimation.
+The primary motivation for employing these two models is to accelerate the evaluation process of a large number of candidate circuits.
+Specifically, we compare the inference time of 100 GNNs-1 predictions for PST against the time required for directly computing PST values, and similarly, we compare the inference time of 100 GNNs-2 predictions for KTA with their direct calculation.
+As illustrated in Fig.
+9 and summarized in Table 1, it is important to note that the y-axis in Fig.
+9 is presented on a logarithmic scale.
+Table 1: Runtime of Prediction and Direct Calculation for PST and KTA
+Metric Prediction ($\cdot s$) Calculation ($\cdot s$) Speedup
+PST $0.0012\pm 0.0000$ $0.4572\pm 0.0845$ $\bm{381\times}$
+KTA $0.0012\pm 0.0000$ $70.0540\pm 22.2887$ $\bm{58378\times}$
+Compared to direct calculation, using GNNs-1 and GNNs-2 for predicting PST and KTA achieves speedups of 381×381$\times$381 × and 58378×58378$\times$58378 ×, respectively.
+These substantial accelerations enable the efficient evaluation of a large number of candidate circuits, making large-scale quantum kernel screening feasible.
+In addition to evaluating the acceleration in computation time, we further investigate the scalability of GNNs-1 and GNNs-2 with respect to the number of qubits.
+Detailed analysis and results on qubit scalability are provided in Appendix B.
+IV.2 Performance of HaQGNN on benchmarks
+Model and Training Setups.
+After evaluating the candidate circuits using GNNs-1 and GNNs-2, we select the top-$k$ circuits that exhibit both high KTA and high PST.
+In our experiments, we set $k=10$, i.e., the top 10 high-performing quantum kernels are selected.
+These quantum kernels are then fine-tuned by training their circuit parameters from scratch.
+The training objective is to maximize classification accuracy using support vector machines (SVMs) based on the corresponding quantum kernels.
+During the parameter training process, we employ the Adam optimizer with a learning rate of 0.01 to efficiently update the trainable parameters.
+We further assess the performance of the proposed method across multiple noisy simulators that incorporate noise models derived from IBM Perth, IBM Lagos, IBM Nairobi, IBM Jakarta, and IBM Torino.
+We compare our proposed method against seven representative baselines:
+Random: The 25 quantum kernels are randomly sampled from the gate set, and the average performance is reported.
+Training Embedding Kernels (TEK) [24]: A method that optimizes parameters on a fixed circuit structure.
+In our setup, each TEK kernel consists of four blocks, where each block contains two layers of parameterized 1-qubit gates followed by a layer of 2-qubit gates connecting all qubits in a ring topology.
+Radial Basis Function Kernel (RBFK) [48, 49]: A widely used classical kernel with the kernel matrix defined as $\left[K\right]_{i,j}=exp\left(-\gamma\left\|x_{i}-x_{j}\right\|^{2}\right)$, where $\gamma$ is a hyperparameter.
+In the simulations, $\gamma$ is set to the reciprocal of the input feature vector dimension, which serves as a reasonable heuristic to balance the sensitivity of the kernel.
+QuantumSupernet [50]: Most hyperparameters follow the settings in [50], with one modification: we replace full-batch gradient descent with mini-batch gradient descent (batch size = 32) to ensure efficient and thorough parameter updates during SuperCircuit training.
+QuantumNAS [8]: We adopt the same hyperparameter settings as in [8] for both SuperCircuit training and evolutionary search.
+The gate set is adapted to match the constraints of each target quantum device.
+EliVagar [7]: This method also incorporates early rejection of low-fidelity circuits.
+We follow the hyperparameter settings provided in [7] and adapt the gate set to match the supported operations of the target quantum devices.
+QuKerNet [23]: A recent method that leverages neural predictor for quantum kernel design.
+In our experiments, we follow the hyperparameter configuration provided in [23] and simulate the circuits with noise information derived from real quantum devices.
+Figure 10: Classification accuracy of different methods on the Credit Card (CC) task under noisy simulation.
+In simulation experiments, if the logical circuits generated by the aforementioned baselines are incompatible with the target quantum hardware, we adopt Qiskit’s default mapping and routing strategies for compilation.
+Specifically, logical qubits are mapped to physical qubits in order, and SWAP gates are heuristically inserted to adapt the circuit to the device topology.
+Dataset Setup.
+We conduct experiments on three different QML benchmark tasks: Credit Card (CC) Fraud Detection (binary classification: fraudulent vs.
+not fraudulent), MNIST-5 (five-class classification: digits 0, 1, 2, 3, and 4), and FMNIST-4 (four-class classification: T-shirt, Trouser, Bag, and Ankle Boot).
+For the MNIST-5 and FMNIST-4 benchmarks, we construct balanced training and test datasets by selecting 2000 training samples and 400 test samples from the original datasets.
+The original 784-dimensional images are reduced to 40 dimensions using the minimum Redundancy Maximum Relevance (mRMR) algorithm [23].
+For the Credit Card (CC) task, we sample 1000 instances from the original dataset while ensuring that at least 5% of the samples are labeled as fraudulent.
+We randomly split 30% of the data as the test set.
+The original 28-dimensional features are reduced to 14 dimensions using mRMR.
+To validate the performance advantage of the quantum kernels discovered by the HaQGNN method, we conduct experiments on a 4-qubit Credit Card (CC) task using two IBM quantum devices with different qubit scales: a 7-qubit device (IBM Perth) and a 133-qubit device (IBM Torino).
+The results are presented in Fig.
+10.
+Our HaQGNN method achieves the highest classification accuracy among all compared approaches, with particularly significant improvements observed on the larger 133-qubit IBM Torino device.
+This performance gain can be attributed to the hardware-aware nature of HaQGNN, which allows the method to generate quantum circuits that are adapted to the physical characteristics of the device.
+On larger quantum processors, HaQGNN is capable of selecting qubits that are less affected by noise and synthesizing high-fidelity circuits accordingly, thereby leading to superior performance in real-device executions.
+Furthermore, we evaluate the 7-qubit MNIST-5 classification task on five IBM quantum backends: IBM Perth, IBM Lagos, IBM Nairobi, IBM Jakarta, and IBM Torino.
+As shown in Fig.
+11, the proposed HaQGNN method consistently achieves superior classification accuracy compared to other baselines.
+On IBM Lagos and IBM Jakarta, the classification performance of HaQGNN is slightly lower than that of QuantumNAS, EliVagar, and QuKerNet, respectively.
+This is primarily because these devices only have 7 qubits, which limits the capability of hardware-aware design to perform effective qubit selection—many qubits are significantly affected by noise, reducing the advantage of noise-aware circuit generation.
+However, when evaluated on the 133-qubit IBM Torino simulator, HaQGNN demonstrates a clear performance advantage, achieving the highest classification accuracy among all methods due to its ability to exploit lower-noise qubits through hardware-aware circuit design.
+Since the FMNIST-4 benchmark task involves 8 qubits, simulations were conducted exclusively on the IBM Torino device.
+As shown in Fig.
+12, the HaQGNN method achieves a significantly higher classification accuracy compared to other methods.
+Figure 11: Classification accuracy of different methods on the MNIST-5 classification task under noisy simulation.
+Figure 12: Classification accuracy of different methods on the FMNIST-4 classification task under noisy simulation.
+V Conclusion
+In this work, we introduced HaQGNN, a hardware-aware quantum kernel design method that leverages graph neural networks to evaluate and select task-specific quantum circuits under the constraints of NISQ hardware.
+By incorporating real-device topology, noise characteristics, and feature selection into the design pipeline, HaQGNN enables scalable and efficient kernel construction tailored for near-term quantum devices.
+The use of two models (GNNs-1 and GNNs-2) allows for fast and accurate prediction of circuit fidelity and kernel performance, significantly reducing the computational cost of kernel screening.
+Experimental results on three classification benchmarks, Credit Card (CC), MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms state-of-the-art quantum kernel methods in terms of accuracy.
+These findings highlight the potential of learning-based, hardware-aware approaches for bridging the gap between theoretical quantum kernel methods and practical deployment on noisy quantum devices.
+As a future direction, exploring the transferability and generalization capabilities of quantum kernels across different tasks or data distributions may further enhance the practicality and reusability of quantum feature maps in broader quantum machine learning applications.
+Acknowledgements.
+This work is supported by the National Natural Science Foundation of China (Grant No.
+62471126), the Jiangsu Key R&D Program Project (Grant No.
+BE2023011-2), and the Fundamental Research Funds for the Central Universities (Grant No.
+2242022k60001).
+Appendix A Strategy for selecting subgraphs from the device topology
+In our investigation of quantum device noise characteristics, we observed a rapid decay in certain types of errors for large-scale quantum devices, such as the 133-qubit IBM Torino.
+Specifically, the 1-qubit gate error, 2-qubit gate error, and readout assignment error exhibit a sharp decline when ranked across qubits.
+As illustrated in Fig.
+13, (a) presents the $R_{x}$ gate errors sorted in descending order, while (b) shows the readout assignment errors also sorted from highest to lowest.
+Algorithm 2: Select subgraph from the device topology
+1:Device topology $G^{\ast}\left(\left\{V_{i}^{\ast}\right\},\left\{E_{j}^{\ast}\right\}\right)$; Excluded value $E_{xc}$; The number of qubits $N$; Noise types $T_{noise}$.
+2:Subgraph $G$.
+3:for $type$ in range($T_{noise}$do
+4:     if $type$ related to a single qubit then
+5:         Sort all qubits by error in descending order and select the top $E_{xc}$ qubits with the highest error, denoting their associated vertices and edges as $\left\{V^{\ast}\right\}$ and $\left\{E^{\ast}\right\}$.
+6:         $G^{\ast}=G^{\ast}\setminus\left(\left\{V^{\ast}\right\},\left\{E^{\ast}\right\}\right)$triangleright Remove $\left\{V^{\ast}\right\}$ and $\left\{E^{\ast}\right\}$ from topology $G^{\ast}$.
+7:     else if $type$ related to two qubits then
+8:         Sort all connections by error in descending order and select the top $E_{xc}$ connections with the highest error, denoting their associated edges as $\left\{E^{\ast}\right\}$.
+9:         $G^{\ast}=G^{\ast}\setminus\left\{E^{\ast}\right\}$triangleright Remove $\left\{E^{\ast}\right\}$ from topology $G^{\ast}$.
+10:     end if
+11:end for
+12:Select a subgraph $G$ from the topology $G^{\ast}$ with $N$ qubits and high degree of connectivity. $\triangleright$ The degree of connectivity is defined as the maximum vertex degree.
+13:return Subgraph $G$
+From Fig.
+13 (a), we observe that after excluding the top 14 qubits with the highest $R_{x}$ gate errors (highlighted by blue bars), the remaining 119 qubits exhibit relatively uniform $R_{x}$ errors.
+A similar trend is observed in Fig.
+13 (b): after excluding the top 18 qubits with the highest readout errors, the error differences among the remaining qubits become much smaller.
+Based on these observations, we propose a subgraph selection algorithm to identify qubits with more consistent noise characteristics.
+The detailed procedure is presented in Algorithm A.
+The main idea of Algorithm A is to remove vertices (qubits) and connecting edges that are heavily affected by noise from the device topology, resulting in a subgraph with low noise impact.
+(a) $R_{x}$ gate errors of different qubits on IBM Torino (sorted in descending order)
+(b) Readout assignment errors of different qubits on IBM Torino (sorted in descending order)
+Figure 13: Noise characteristics of different qubits on IBM Torino.
+Appendix B Qubit scalability of GNNs-1 and GNNs-2
+To investigate the qubit scalability of GNNs-1 and GNNs-2, we conducted a simulation experiment based on the IBM Torino device, which features 133 qubits.
+Specifically, we extended the “Target Qubit” component of the node feature vectors used in Fig.
+5 from the dimension of 16 to 21 to accommodate larger subgraphs.
+Fig.
+14 presents the scalability performance of GNNs-1 for predicting PST and GNNs-2 for predicting KTA across varying qubit numbers.
+Figure 14: Qubit scalability of GNNs-1 and GNNs-2 for PST and KTA prediction.
+Overall, we observe that as the number of qubits increases, models trained on smaller subgraphs experience notable performance degradation when applied to circuits with more qubits.
+In particular, models trained on $N=4$ qubits show a sharp decline in predictive accuracy as qubit count increases.
+In contrast, models trained on larger subgraphs (e.g., $N=7$ or $8$) exhibit only a gradual decline in performance, suggesting improved scalability.
+These findings imply the existence of a threshold in model generalizability across qubit counts, aligning with the results reported in [51], where convolutional neural networks (CNNs) were used to predict expectation values of quantum circuits.